OpenStudy (anonymous):

write the equation of a line with slope of -2 which contains the point 3,6.. Show all your work !!

OpenStudy (niel):

mark point 3,6 and go down 2 and right 1 and mark it... then contoinue the line

OpenStudy (anonymous):

General formula for linear equation : y = ax+b Given slope is -2 : y = -2x+ b To find b, solve the eqn : 6 = -2(3) + b 6 + 6 =b b= 12. End formula : y = -2x + 12

OpenStudy (anonymous):

Both method are correct, altrough I express a frank bias for algebraic only proceding.

OpenStudy (niel):

lol true...